OpenStudy (anonymous):

Use the value of the trigonometric function to evaluate the indicated functions such as given cos(t) = 3/5. 1) cos( pi - t) 2) cos ( t + pi)

OpenStudy (anonymous):

so cos (t) = -(-3/5)

OpenStudy (anonymous):

how do you apply the odd even formulas to the problem?

OpenStudy (anonymous):

cosine and secant are even but i'm not understanding how the answer changes to equal negative 3/5 form positive 3/5

OpenStudy (er.mohd.amir):

or u can open cos (A+B)=cos A cos B -sin A sin B and cos (A-B)=cos A cos B + sin A sin B

OpenStudy (anonymous):

I'm confused how odd and even functions work solving this problem

OpenStudy (er.mohd.amir):

what are even and odd function,tell me thhen i expalin u.

OpenStudy (anonymous):

says given sin t = 4/5 then problem sin(pi-t) = 4/5 and sin(t+ pi) = -4/5 ...sin is an odd function

OpenStudy (er.mohd.amir):

sorry

OpenStudy (er.mohd.amir):

i write formule for pi/2

OpenStudy (anonymous):

cos and sec are the only even trig functions

OpenStudy (er.mohd.amir):

cos (pi-x)=-cos x cos (pi+x)=-cos x

OpenStudy (anonymous):

can you explain how that equals -cos x

OpenStudy (anonymous):

is this because cos is the x value which never changes?

OpenStudy (er.mohd.amir):

since (pi-x) is in II Quradent and in II cos is -ve and (pi+x) is in III and cos is -ve in III cos is +ve only in I and IV quradent.

OpenStudy (anonymous):

so knowing the formula how can i write one of those two problems would this be correct cos(pi - x) = cos(pi -(3/5)) = cos(pi) -cos(3/5) so cos = -3/5

OpenStudy (er.mohd.amir):

No cos (pi-x)=-cos x=-3/5

OpenStudy (anonymous):

so cos(x+pi) = -cos x = -3/5

OpenStudy (er.mohd.amir):

understand or not

OpenStudy (anonymous):

not really.. im looking at an easier problem Given: sin(-t) = 4/7 So this can be stated as sin(-t) = - sin(t) = -(-4/7) .... why is this not -4/7

OpenStudy (er.mohd.amir):

if sin(-x)=a then -sin x=a or sin x=-a

OpenStudy (anonymous):

so if cos x = -3/4 given cos(-t) = -3/4 sec(-t) = -4/3

OpenStudy (mathmale):

I think your best bet is to use the difference formula for the cosine function to answer this question. cos (a-b) = cos a cos b + sin a sin b. Please evaluate cos (pi - t). Think carefully: what is the value of cos pi?